OpenStudy (anonymous):

3 sisters are each 2 years apart in age. The youngest sister is Emily, the middle sister is Samantha and the oldest sister is Lauren. Together their combined ages are 42. How old is each sister? a. Define the variable(s). b. Write an equation for this situation using one variable. c. Solve the equation for the ages of the 3 sisters.

OpenStudy (anonymous):

@texaschic101

OpenStudy (vishweshshrimali5):

Did you give it a try ?

OpenStudy (anonymous):

no

OpenStudy (vishweshshrimali5):

Ohk then let me give a hint

OpenStudy (vishweshshrimali5):

Let the age of: Emily = x Samantha = y Lauren = z

OpenStudy (anonymous):

guage! its easy haha

OpenStudy (vishweshshrimali5):

Since, the difference between the ages is 2 years then, can we write this? y = x+2; z = y+2

OpenStudy (vishweshshrimali5):

Thus: y = x+2 ---- (1) z = (x+2)+2 = x + 4 ------ (2) Also, you know the combined sum of the ages of the three sisters, so: x + y + z = 42 --- (3) Now, use equations (1) and (2) to write equation (3) in terms of x only

OpenStudy (anonymous):

kyle my name is spelled gaige

OpenStudy (anonymous):

@kylefoster58 and your sitting next to me lol

OpenStudy (vishweshshrimali5):

@mathsciencehistory can you write eq (3) in terms of x only ?

OpenStudy (anonymous):

lol maybe

OpenStudy (anonymous):

x+x+x=42

OpenStudy (vishweshshrimali5):

Nope! Use equations (1) and (2) to replace y and z by x + 2 and x+4 respectively

OpenStudy (anonymous):

no i spelled it right!

OpenStudy (anonymous):

lmao kial

OpenStudy (texaschic101):

variables are E, S, L 2 years apart S = E + 2 L = E + 4 E + S + L = 42 E + (E + 2) + (E + 4) = 42 3E + 6 = 42 3E = 42 - 6 3E = 36 E = 12 now sub 12 in for E in the other equations to find their ages

OpenStudy (anonymous):

im soo lost i learn better if you just write out the answer

OpenStudy (texaschic101):

if E = 12...then.. S = E + 2 S = 12 + 2 S = 14 and L = E + 4 L = 12 + 4 L = 16 so emily is 12, samantha is 14, and lauren is 16

OpenStudy (anonymous):

ok thanxs

OpenStudy (texaschic101):

sure thing :)

OpenStudy (anonymous):

and what would the equation with one variable be??

OpenStudy (texaschic101):

E + (E + 2) + (E + 4) = 42

OpenStudy (texaschic101):

thats if you let E represent Emily

OpenStudy (texaschic101):

understand ?

OpenStudy (anonymous):

sorta ill come back to these later when im done with the exam